OpenStudy (anonymous):

Let the limit as x approaches -9; f(x)=9. Find limit as x approaches -9; square root of f(x)

OpenStudy (anonymous):

yes, thats the problem.

OpenStudy (agent0smith):

No, I mean, use this to find the answer...\[\huge \lim_{x \rightarrow a} \sqrt {f(x) } = \sqrt {\lim_{x \rightarrow a} f(x) }\]

OpenStudy (anonymous):

that's exactly how the problem looks still confused

OpenStudy (anonymous):

so then it should come as 3 ?

OpenStudy (anonymous):

thanks

OpenStudy (agent0smith):

You know what \[\huge \lim_{x \rightarrow -9} f(x) = 9\] and this\[\huge \lim_{x \rightarrow -9} \sqrt {f(x) } = \sqrt {\lim_{x \rightarrow -9} f(x) }\]

OpenStudy (agent0smith):

Understand?

OpenStudy (anonymous):

yes thanks man!